# An Exact Solution Algorithm for the Bi-Level Optimization Problem of Electric Vehicles Charging Station Placement

### v1

@@ 7 Conclusion
Our methodology represents an important advancement over existing approaches in the charging station location literature, which resulted in both superior solution quality and computational efficiency. While existing heuristic methods provide no optimality guarantees and often struggle with solution quality, our algorithm consistently produces exact optimal solutions with mathematical certainty. More critically, compared to exact solution methods from the literature [doi:10.1287/trsc.2021.0494, 10122631, 10034431, 9690622] that require hundreds of hours of computational time to solve smaller network instances, often without achieving optimality gaps below 1%, our algorithm demonstrates considerable efficiency by terminating within minutes while guaranteeing global optimality. This computational superiority, spanning over two orders of magnitude in runtime reduction, transforms charging infrastructure planning from a computationally prohibitive problem requiring days of computation into a tractable optimization task that can be solved in real time. Specifically, our computational experiments demonstrate rapid convergence, minimal optimality gaps, and substantially reduced runtimes. The algorithm will effectively handle extensive scenarios involving thousands of candidate combinations. Our results indicate an outstanding computational and theoretical improvement over previous approaches in the literature.
Future research can consider adding dynamic operational constraints, such as real time station congestion, queuing effects, and variable charging rates based on location. Furthermore, evaluating uncertainties in user behavior, demand forecasts, and technological advancements in battery capacities can enhance practical applicability. Additionally, the algorithm runtime can improved by using faster algorithms like traffic assignment by paired alternative segments [BARGERA20101022] instead of Frank Wolfe that we implement.

### v2

@@ 8 Conclusion
Our methodology represents an important advancement over existing approaches in the charging station location literature, which resulted in both superior solution quality and computational efficiency. While existing heuristic methods provide no optimality guarantees and often struggle with solution quality, our algorithm consistently produces exact optimal solutions with mathematical certainty. More critically, compared to exact solution methods from the literature [13, 27, 20, 19] that require hundreds of hours of computational time to solve smaller network instances, often without achieving optimality gaps below 1%, our algorithm demonstrates considerable efficiency by terminating within minutes while guaranteeing global optimality. Compared to existing methods in the literature, our algorithm solves networks one to two orders of magnitude larger while achieving tighter optimality gaps in comparable or shorter runtimes, transforming charging infrastructure planning from a computationally prohibitive problem into a tractable optimization task. Specifically, our computational experiments demonstrate rapid convergence, minimal optimality gaps, and substantially reduced runtimes. The algorithm will effectively handle extensive scenarios involving thousands of candidate combinations. Our results indicate an outstanding computational and theoretical improvement over previous approaches in the literature.
Future research can extend the congestion model to capture real-time dynamic queuing, such as real time station congestion. Furthermore, evaluating uncertainties in user behavior, demand forecasts, and technological advancements in battery capacities can enhance practical applicability. Recent work by Rashgi Shishvan et al. [21] shows that the spatial distribution of charging infrastructure considerably affects load profile smoothness and ramping needs, which suggests that coupling transportation network optimization with power system objectives could yield more holistic infrastructure planning. Additionally, the algorithm runtime can be improved by using faster algorithms like traffic assignment by paired alternative segments [2] instead of Frank Wolfe that we implement.